OpenStudy (anonymous):

Divide 32x^4 - 12x^2 + 7x by 2x .

OpenStudy (anonymous):

Just divide every term by 2x

OpenStudy (anonymous):

obviously. but does anything happen to the exponents or the x?

OpenStudy (anonymous):

\[\frac{ 2*16*x*x*x*x-2*6*x*x+2*3.5*x }{ 2x }\]

OpenStudy (anonymous):

Does that make it simpler?

OpenStudy (anonymous):

no im sorry can you explain it in words instead of numbers for please?

OpenStudy (anonymous):

Take off 2x from every term and what's left is your answer. I laid it out for you very well.

OpenStudy (anonymous):

they decrease by 1

OpenStudy (anonymous):

\(\frac{x^n}{x}=x^{n-1}\)

OpenStudy (anonymous):

a\[a^{24}-a ^{18}+a ^{12} \div a^{6}\]